OpenStudy (anonymous):

It takes one crew 6 hours to clean an auditorium after an event. If a second crew helps, it only takes 2 hours. How long would it take the second crew, working alone, to clean the auditorium? @Algebra Help!

OpenStudy (anonymous):

first crew has rate of \[\frac{1}{6}\] put second crews rate as R then \[(\frac{1}{6}+R)\times 2=1\] solve for R

OpenStudy (anonymous):

get \[\frac{1}{6}+R=\frac{1}{2}\] \[R=\frac{1}{2}-\frac{1}{6}=\frac{1}{3}\] to there rate is 1/3 per hour and it would take them 3 hours working alone